OpenStudy (anonymous):

help please?? how do i write -3i in trigonometric form?

OpenStudy (tkhunny):



OpenStudy (michele_laino):

the complex number \(-i\) is this:  so we have to consider the phase \(\phi=3 \pi/2\)

OpenStudy (michele_laino):

next, please apply the Euler formula: \[\large {e^{i\phi }} = \cos \phi + i\sin \phi \]

OpenStudy (mathmale):

In other words, starting from the positive real axis, we travel in a circle counterclockwise, passing the positive imaginary axis and the negative real axis, to arrive at the negative imaginary axis. This counterclockwise motion is described by the angle 3pi/2. To write -3i in trig form, you'll need to imagine or actually draw a circle of radius ( what? ). Last step: Describe this destination point in the form (r, theta). r is the radius, theta is the angle. Try it, please.

OpenStudy (michele_laino):

please replace \(\phi=3 \pi/2\) into both sides of the Euler formula